Circular No. 7289

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 T3 (LINEAR) Linkage at the Minor Planet Center of observations by LINEAR (dis-

covery observation below) on several nights during Oct. 3–21 revealed an object with a nearly-parabolic retrograde orbit. This orbit also represented a single-night detection of an object by E. W. Elst and S. Ipatov at Uccle on Oct. 18. Following placement of an ephemeris on The NEO Confirmation Page further observations were made on Oct. 24 and 25. In response to enquiries, Elst remarked that the object was diffuse and had a possible tail to the north; J. Tichá and M. Tichý, Kleť, reported the object as slightly more diffuse than stars of comparable brightness and deduced a coma size of 9′′; and D. Durig, Sewanee, TN, in poor conditions (strong wind, full moon), also noted the object’s diffuse appearance.

1999 UT  α_{2000}  δ_{2000}  m_{2}
Oct. 3.33731  3^{h}09^{m}03.^{s}31  +22^{o}06′48′′.3  18.3

Astrometry and the following parabolic orbital elements appear on MPEC 1999-U27:

T = 2000 Aug. 29.1385 TT  ω = 210.^{o}9351
Ω = 223.5095 }2000.0
q = 5.370459 AU  i = 104.6325
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 19 2^{h}58.^{m}44 +19^{o}46.^{′}6 4.920 5.851 157.^{o}1 3.^{o}8 17.1
29 2 50.74 +18 06.5 4.843 5.822 169.3 1.8 17.1
Nov. 8 2 42.77 +16 21.1 4.804 5.795 178.1 0.3 17.0
18 2 34.92 +14 33.7 4.804 5.768 165.8 2.4 17.0
28 2 27.60 +12 48.1 4.843 5.742 153.4 4.4 17.0
Dec. 8 2 21.13 +11 07.8 4.915 5.716 141.2 6.2 17.0
18 2 15.78 + 9 35.5 5.018 5.692 129.2 7.7 17.1
28 2 11.68 + 8 13.0 5.144 5.668 117.6 8.8 17.1

NOVA IN IC 1613 M. M. M. Santangelo reports that prediscovery CCD frames (limiting

mag R ∼ 20.0), taken on Sept. 11.09 UT with the Monte Agliale 0.51m reflector, show the nova in IC 1613 (cf. IAUC 7287) at R = 19.1; his astrometry yields α = 1^{h}04^{m}49^{s}.04, δ = +2^{o}05′29′′.1 (equinox 2000.0).

1999 October 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green